# Learning to Bid in Non-Stationary Repeated First-Price Auctions

### v2

@@ 1.1 Our Contributions
  • We propose policies that are efficiently implementable and achieve the minimax-optimal dynamic regret guarantees of $\tilde{O}\left(\sqrt{TV_{T}}\right)$ and $\tilde{O}\left(L_{T}\right)$, where $\tilde{O}(\cdot)$ hides poly-logarithmic factors. The one-sided Lipschitzness of the reward function poses significant challenges in predicting the optimal bid, as discussed further in Section 4.1.1. To address this challenge, we employ the Optimistic Mirror Descent (OMD) framework (Chiang et al. 2012, Rakhlin and Sridharan 2013), a powerful tool with a customizable optimism vector that achieves improved static regret rates in slowly evolving environments. Interestingly, we design the optimism not simply to minimize the static regret, but rather to achieve a favorable balance between the static regret and the transition cost from static regret to dynamic regret, thus achieving the optimal dynamic regret rates.
@@
    Table 1: Dynamic regret rates lower bounds and upper bounds when the either $V_{T}$ or $L_{T}$ is constrained. $V_{T}\coloneqq\sum_{t=2}^{T}|m_{t}-m_{t-1}|$ and $L_{T}\coloneqq\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1})$ are two metrics to measure the regularity of the opponents’ highest bid sequence. Here we use $\tilde{O}(\cdot)$ to omit polylogarithmic factors.
@@ 1.4 Notations
Additionally, $\mathbbm{1}(\cdot)$ denotes the indicator function of an event. $\mathbb{E}[\cdot]$ represents the expectation operator. $[s]\coloneqq\{1,\dots,s\}$ denotes the set of integers from 1 to $s$. For a convex and differentiable function $\psi$ defined on a convex region $\mathcal{P}$, $D_{\psi}(p,q)\coloneqq\psi(p)-\psi(q)-\langle p-q,\nabla\psi(q)\rangle$ is the Bregman divergence. We use $\mathbf{1}$ to denote an all-ones vector. We use standard asymptotic notation $O(\cdot)$, $\Omega(\cdot)$, $\Theta(\cdot)$ and $\tilde{O}(\cdot)$ to simplify the analysis: We use $x_{n}=O(y_{n})$ to denote that there exist constants $n_{0}\in\mathbb{N}^{+}$ and $M\in\mathbb{R}^{+}$ such that for all $n\geq n_{0}$, $x_{n}\leq M\cdot y_{n}$. Similarly, $x_{n}=\Omega(y_{n})$ is equivalent to $y_{n}=O(x_{n})$, $x_{n}=\Theta(y_{n})$ means $x_{n}=O(y_{n})$ and $x_{n}=\Omega(y_{n})$, and $\tilde{O}(\cdot)$ is similar to $O(\cdot)$ but hides polylogarithmic factors.
@@ 3.1 Problem Formulation
Here, $\mathbbm{1}(b_{t}\geq m_{t})$ is the indicator function that equals $1$ if the bidder wins the auction (i.e., $b_{t}\geq m_{t}$), and 00 otherwise. The one-sided Lipschitz property means: when moving from a higher bid $b^{\prime}\leq v_{t}$ to a slightly lower bid $b$, the reward increase is bounded by the bid difference, but the reward decrease can be significant (particularly when crossing the discontinuity at $m_{t}$).
For simplicity, we assume the time horizon $T$ is known to the learner. If $T$ is unknown, the doubling trick (Auer et al. 2002, Cesa-Bianchi and Lugosi 2006) can be used to eliminate this requirement. Since this is a sequential decision-making problem, it is essential to formally define the information received by the learner before submitting $b_{t}$. We mainly consider the case where the learner observes $m_{t}$, the highest bid from other bidders, so the information up to time $t-1$ can be described by the following filtration:
@@
where $\sigma(\cdot)$ is the $\sigma$-algebra generated by the observations. Conventionally, the filtration up to $t-1$ should not include $v_{t}$, as this represents information from the current round. However, prior work (Han et al. 2020, Balseiro et al. 2023, Han et al. 2025) assumes that the bidder knows $v_{t}$ before determining their bid $b_{t}$. This assumption is reasonable because ad exchanges typically display the ad impression and related contextual or demographic information to bidders, enabling them to estimate the value of the impression. Therefore, we include $v_{t}$ in the filtration.
@@ 3.2 The Optimistic Mirror Descent Framework
$a_{t,i}\leftarrow\begin{cases}0,\quad&(\text{Option \text{I}})\\ 4\eta(r_{t,i}-\mu_{t})^{2},\quad&(\text{Option \text{II}})\\ \end{cases}$
@@ 4.1.1   Why Existing Works Do Not Directly Apply?
$\begin{split}&\mathbb{E}[\text{DR}_{T}(\pi)]=\sup_{b_{1}^{*},\dots,b_{T}^{*}\in[0,1]}\sum_{t=1}^{T}\left(r(b_{t}^{*};v_{t},m_{t})-\mathbb{E}[r(b_{t};v_{t},m_{t})]\right)\\ =&\sum_{j=1}^{n}\left(\max_{f\in\tilde{\mathcal{F}}}\sum_{t\in\mathcal{T}_{j}}r(f(v_{t});v_{t},m_{t})-\sum_{t\in\mathcal{T}_{j}}\mathbb{E}[r(b_{t};v_{t},m_{t})]\right)\\ &+\sum_{j=1}^{n}\left(\sum_{t\in\mathcal{T}_{j}}r(b_{t}^{*};v_{t},m_{t})-\max_{f\in\tilde{\mathcal{F}}}\sum_{t\in\mathcal{T}_{j}}r(f(v_{t});v_{t},m_{t})\right)\\ \coloneqq&\sum_{j=1}^{n}\mathcal{S}^{\mathcal{A}}(\tilde{\mathcal{F}},\mathcal{T}_{j})+\sum_{j=1}^{n}\mathcal{C}(\tilde{\mathcal{F}},\mathcal{T}_{j}).\\ \end{split}$ (6)
@@
To demonstrate the application of the decomposition in Equation (6) for achieving sublinear dynamic regret, we consider using the restart scheme with the Hedge algorithm as $\mathcal{A}$. We partition the time horizon into batches of equal length $\Delta_{T}$, with the possible exception of the last batch. Let $\tilde{\mathcal{F}}$ be the set of constant policies, i.e., ~$\tilde{\mathcal{F}}\coloneqq\{f(v;\tau)=\tau|\tau\in[0,1]\}$. Then, the following proposition holds:
@@
where we replace the minimax-optimal policy $\mathcal{A}$ and class $\tilde{\mathcal{F}}$ with a potentially different policy $\mathcal{A}^{\prime}$ and class $\mathcal{F}^{\prime}$, aiming for a regret guarantee that scales with the intra-batch temporal variation. The step marked with ${{\leq}}$ is the crux of our approach and requires establishing that the static regret can indeed scale with $V_{T,j}$ within each batch (to be elaborated in the sequel).
@@
$j\leftarrow 1,\eta\leftarrow\frac{1}{2},\epsilon\leftarrow\frac{1}{T},c\leftarrow\frac{1}{T}$;
@@ 4.2 Dynamic Regret Rates under the Switching Number Constraint
$j\leftarrow 1$,$t\leftarrow 1$;
while $t\leq T$ do
@@ 4.3 Best-of-Both-Worlds Dynamic Regret
$j\leftarrow 1$,$t\leftarrow 1$;
for $t\leftarrow 1$ to $T$ do
@@ 5 Dynamic Regret Lower Bounds
In this section, we demonstrate how to establish minimax lower bounds for the class of auction instances where the opponents’ highest bid sequence is constrained by either $V_{T}$ or $L_{T}$. Our main effort focuses on the case of $V_{T}$. Following Besbes et al. (2015), we term $V_{T}$ the variation budget. For the corresponding lower bound construction, we partition the time horizon $T$ into batches of equal size $H$, and we allocate a small and fixed amount of variation budget $\frac{1}{H}$ to each batch to create a jump at locations drawn from the uniform distribution. Within each batch, due to the one-sided Lipschitzness of the reward function, the learner faces a dilemma: providing a small bid incurs 00 dynamic regret before the jump occurs, but will incur $\Omega(1)$ dynamic regret at the jump point. Bidding a higher price avoids the $\Omega(1)$ dynamic regret at the jump point, but incurs $\frac{1}{H}$ dynamic regret for each round until the jump point. Formally, we show that any non-anticipatory policy incurs $\Omega(1)$ dynamic regret within each batch based on dynamic programming in Lemma 5.1. Since there are $\Theta(T/H)$ batches, choosing $H=\Theta(\sqrt{T/V_{T}})$ satisfies the variation budget constraint and also implies the $\Omega(\sqrt{TV_{T}})$ dynamic regret lower bound. For the case of $L_{T}$, we achieve the $\Omega(L_{T})$ lower bound by reducing to the case of $V_{T}$, which is possible because each batch considered in Lemma 5.1 contains only one random jump.
@@
Figure 1: An illustration of the construction for the lower bound. In odd-numbered batches, $m_{t}$ jumps from 00 to $\frac{1}{H}$ at random locations, while in even-numbered batches, $m_{t}$ jumps from $\frac{1}{H}$ back to 00. The parameter $H$ is carefully chosen to satisfy the variation budget constraint and ensure that we obtain the desired lower bound.
@@ 6.2 Performance Against Budget-Pacing Bidders
Robustness analysis: To ensure that our advantage arises from algorithmic adaptivity rather than simply from exploiting slowly varying sequences, we conducted additional experiments where both the learner’s and the opponents’ values are drawn from the same distributions: uniform, a truncated Gaussian (with mean $0.4$ and standard deviation $0.2$), or Beta(3,3)33(3,3)( 3 , 3 ). In these experiments, we set $T=12000$ with $20$ budget-pacing bidders and vary the initial budget of each bidder. We repeat each experiment $50$ times and report the average performance. As shown in Figure 5, our algorithms remain competitive in regimes with sufficient budgets and outperform baselines in regimes with insufficient budgets, confirming that their adaptive capabilities extend beyond merely capitalizing on slow variation.
@@ 7 Conclusion
We thank Yanjun Han from New York University for providing the code of the SEW policy.

### v3

@@ 1.1 Our Contributions
[…]
@@
[…]
@@ 1.4 Notations
[…]
@@ 3.1 Problem Formulation
Here, $\mathbbm{1}(b_{t}\geq m_{t})$ is the indicator function that equals $1$ if the bidder wins the auction (i.e., $b_{t}\geq m_{t}$), and 0 otherwise. The one-sided Lipschitz property means: when moving from a higher bid $b^{\prime}\leq v_{t}$ to a slightly lower bid $b$, the reward increase is bounded by the bid difference, but the reward decrease can be significant (particularly when crossing the discontinuity at $m_{t}$).
For simplicity, we assume the time horizon $T$ is known to the learner. If $T$ is unknown, the doubling trick (Auer et al. 2002, Cesa-Bianchi and Lugosi 2006) can be used to eliminate this requirement. Since this is a sequential decision-making problem, it is essential to formally define the information received by the learner before submitting $b_{t}$. We mainly consider the case where the learner observes $m_{t}$, the highest bid from other bidders, so the information up to time t1t-1 can be described by the following filtration:
@@
where $\sigma(\cdot)$ is the $\sigma$-algebra generated by the observations. Conventionally, the filtration up to t1t-1 should not include $v_{t}$, as this represents information from the current round. However, prior work (Han et al. 2020, Balseiro et al. 2023, Han et al. 2025) assumes that the bidder knows $v_{t}$ before determining their bid $b_{t}$. This assumption is reasonable because ad exchanges typically display the ad impression and related contextual or demographic information to bidders, enabling them to estimate the value of the impression. Therefore, we include $v_{t}$ in the filtration.
@@ 3.2 The Optimistic Mirror Descent Framework
[…]
@@ 4.1.1  Why Existing Works Do Not Directly Apply?
[…]
@@
To demonstrate the application of the decomposition in Equation (6) for achieving sublinear dynamic regret, we consider using the restart scheme with the Hedge algorithm as $\mathcal{A}$. We partition the time horizon into batches of equal length $\Delta_{T}$, with the possible exception of the last batch. Let $\tilde{\mathcal{F}}$ be the set of constant policies, i.e., $\tilde{\mathcal{F}}\coloneqq\{f(v;\tau)=\tau|\tau\in[0,1]\}$. Then, the following proposition holds:
@@
[…]
@@
[…]
@@ 4.2 Dynamic Regret Rates under the Switching Number Constraint
$j\leftarrow 1$$t\leftarrow 1$;
1exwhile $t\leq T$ do
@@ 4.3 Best-of-Both-Worlds Dynamic Regret
$j\leftarrow 1$$t\leftarrow 1$;
1exfor $t\leftarrow 1$ to $T$ do
@@ 5 Dynamic Regret Lower Bounds
In this section, we demonstrate how to establish minimax lower bounds for the class of auction instances where the opponents’ highest bid sequence is constrained by either $V_{T}$ or $L_{T}$. Our main effort focuses on the case of $V_{T}$. Following Besbes et al. (2015), we term $V_{T}$ the variation budget. For the corresponding lower bound construction, we partition the time horizon $T$ into batches of equal size $H$, and we allocate a small and fixed amount of variation budget $\frac{1}{H}$ to each batch to create a jump at locations drawn from the uniform distribution. Within each batch, due to the one-sided Lipschitzness of the reward function, the learner faces a dilemma: providing a small bid incurs 0 dynamic regret before the jump occurs, but will incur $\Omega(1)$ dynamic regret at the jump point. Bidding a higher price avoids the $\Omega(1)$ dynamic regret at the jump point, but incurs $\frac{1}{H}$ dynamic regret for each round until the jump point. Formally, we show that any non-anticipatory policy incurs $\Omega(1)$ dynamic regret within each batch based on dynamic programming in Lemma 5.1. Since there are $\Theta(T/H)$ batches, choosing $H=\Theta(\sqrt{T/V_{T}})$ satisfies the variation budget constraint and also implies the $\Omega(\sqrt{TV_{T}})$ dynamic regret lower bound. For the case of $L_{T}$, we achieve the $\Omega(L_{T})$ lower bound by reducing to the case of $V_{T}$, which is possible because each batch considered in Lemma 5.1 contains only one random jump.
@@
Figure 1: An illustration of the construction for the lower bound. In odd-numbered batches, $m_{t}$ jumps from 0 to $\frac{1}{H}$ at random locations, while in even-numbered batches, $m_{t}$ jumps from $\frac{1}{H}$ back to 0. The parameter $H$ is carefully chosen to satisfy the variation budget constraint and ensure that we obtain the desired lower bound.
@@ 6.2 Performance Against Budget-Pacing Bidders
Robustness analysis: To ensure that our advantage arises from algorithmic adaptivity rather than simply from exploiting slowly varying sequences, we conducted additional experiments where both the learner’s and the opponents’ values are drawn from the same distributions: uniform, a truncated Gaussian (with mean $0.4$ and standard deviation $0.2$), or Beta(3,3)(3,3). In these experiments, we set $T=12000$ with $20$ budget-pacing bidders and vary the initial budget of each bidder. We repeat each experiment $50$ times and report the average performance. As shown in Figure 5, our algorithms remain competitive in regimes with sufficient budgets and outperform baselines in regimes with insufficient budgets, confirming that their adaptive capabilities extend beyond merely capitalizing on slow variation.
@@ 7 Conclusion
We thank Yanjun Han from New York University for providing the code of the SEW policy. This research is generously supported by the NSF grant CCF-2106508.